a body is projected is at same height from the ground at t=1 s and t=9s. find the max. height attained by it.(g=10m/s)

Grade:9

1 Answers

Sheetal kumari
6 years ago
If t1 and t2 are the time ,when the body is at the same height then, there is a formula from which we find the max height h=1/2×g×t1×t2=1/2×g×1×9(g=10) h=1/2×10×9=45mThus max height reached by it are 45mPlease kindly approved my answer